OpenStudy (anonymous):

a university theater sold 556 tickets for a play, tickets cost $22 per adult tickets and $ 12 for senior citizen ticket, if the total receipt was 8492, how many senior citizen ticket were sold

OpenStudy (kropot72):

Let a = number of adult tickets and s = number of senior tickets sold a + s = 556 22a + 12s = 8492 Multiply the first equation by 22 and get: 22a + 22s = 12232 Subracting equations gives 10s = 3740 s = 374 senior tickets sold

OpenStudy (callisto):

let x and y be the number of senior citizen and adult ticket sold respectively. x+y = 556 => y= 556-x -(1) 12x + 22y = 8492 -(2) Sub (1) into (2) 12x + 22(556-x) = 8492 -10x = 8492-12232 10x = 3740 x=?

OpenStudy (kropot72):

The number of adult tickeys sold is found by the substitution a + 374 = 556 a = 182